# A Framework for Safe Probabilistic Invariance Verification of Stochastic Dynamical Systems

### v1

@@ Abstract
Ensuring safety through set invariance has proven to be a valuable method in various robotics and control applications. This paper introduces a comprehensive framework for the safe probabilistic invariance verification of both discrete- and continuous-time stochastic dynamical systems over an infinite time horizon. The objective is to ascertain the lower and upper bounds of the liveness probability for a given safe set and set of initial states. This probability signifies the likelihood of the system remaining within the safe set indefinitely, starting from the set of initial states. To address this problem, we propose optimizations for verifying safe probabilistic invariance in discrete-time and continuous-time stochastic dynamical systems. These optimizations adapt classical stochastic barrier certificates, which are based on Doob’s non-negative supermartingale inequality, and the equations described in [29, 31], which can precisely define the probability of reaching a target set while avoiding unsafe states. Finally, we demonstrate the effectiveness of these optimizations through several examples using semi-definite programming tools.
@@
Stochastic Systems; Safe Probabilistic Invariance Verification; Liveness Probability; Lower and Upper Bounds.
[…]
@@ 1 Introduction
This paper propose a framework for the verification of safe probabilistic invariance in both stochastic discrete-time and continuous-time systems. The objective of safe probabilistic invariance verification is to compute lower and upper bounds on the liveness probability of the system remaining inside a safe set for all time, starting from a specific initial set. In this framework, several optimizations are proposed for computing these lower and upper bounds. These optimizations are constructed via the adaptation of stochastic barrier certificates in [3, 17], which are built upon the well established Doob’s nonnegative supermartingale inequality [26], and equations in [29, 31], which can precisely characterize the probability of reaching a target set while avoiding unsafe states. To demonstrate and compare the performances of these optimizations, several examples are provided based on the semi-definite programming tool.
@@
    A framework for the safe probabilistic invariance verification in both stochastic discrete-time and continuous-time systems is proposed. The framework not only computes commonly studied lower bounds of the liveness probability, but also provides upper bounds. The incorporation of upper bounds enhances our understanding of the system’s invariance properties and allows for a more accurate estimation of the exact liveness probability.
@@
    The framework incorporates several optimizations to address the safe probabilistic invariance verification. These optimizations are derived from the known nonnegative supermartingale based barrier certificates in [3, 17], as well as the equations proposed by [29, 31] for characterizing the exact probability. It is demonstrated that the optimizations based on barrier certificates are equivalent to those derived from the equations when computing lower bounds. However, for computing upper bounds, the equations from [29, 31] are more effective.
@@
This paper extends our previous conference work [32], which focused solely on the safe probabilistic invariance verification in stochastic discrete-time systems. In contrast to [32], this work expands the method to encompass continuous-time systems. Furthermore, we demonstrate that the optimizations based on known nonnegative supermartingale-based barrier certificates are equivalent to those derived from the equations proposed by [29, 31] when computing lower bounds.
@@ 2.1 Stochastic Discrete-time Systems
The problem addressed in this work is related to the computation of probabilistic invariant sets. These sets define a range of states where a system, starting from any point within the set, must remain within a specified region of interest with a certain probability. By computing probabilistic invariant sets, we can demonstrate that the safe invariance property holds for an initial set if it is a subset of the computed probabilistic invariant sets. Previous works, such as [10], [11], and [8], have approximated polyhedral probabilistic invariant sets using Chebyshev’s inequality for linear systems with Gaussian noise. However, these methods are limited to computing lower bounds of the liveness probability and cannot be applied to compute upper bounds. Furthermore, the above mentioned works have primarily focused on linear systems, whereas this paper considers nonlinear systems.
@@
On the other hand, probabilistic invariance can be evaluated by examining its dual, probabilistic reachability. [2, 1] investigated the finite-time probabilistic invariance problem for discrete-time stochastic (hybrid) systems via reachability analysis. Meanwhile, [24, 25] studied the infinite-time probabilistic invariance by defining it as a finite-time reach-avoid property in combination with infinite-time invariance around absorbing sets over the state space and provided a lower bound for the infinite-time probabilistic invariance. Recently, [7] proposed an algorithm to compute infinite-horizon probabilistic controlled invariant sets which provide a lower bound on the liveness probability based on the dynamic program in [2]. These sets are computed using the stochastic backward reachable set from a robust invariant set. One obstacle faced by the aforementioned methods is the computation of absorbing or robust invariant sets, when they do exist.
@@
Another two closely related works to the present one are [3], which studies the safety and reachability verification based on non-negative supermartingale based barrier certificates, and [29], which formulates a set of equations being able to characterize the exact probability of reaching a specified target set while avoiding unsafe states. With an assumption that the evolution space $\mathcal{X}$ is a robust invariant (i.e., $\bm{f}(\bm{x},\bm{d}):\mathcal{X}\times\mathcal{D}\rightarrow\mathcal{X}$), barrier certificates inspired by the ones in [17] and [18] were formulated for safety and reachability verification of stochastic discrete-time systems in [3]. The first set of optimizations for discrete-time systems in this paper was adapted from them. The assumption that the evolution space $\mathcal{X}$ (i.e., the safe set in this paper) is a robust invariant is abandoned in our method. Instead, an auxiliary switched system with a robust invariant set is borrowed to construct our constraints for addressing the safe probabilistic invariance verification problem. The second set of optimizations proposed in this paper is inspired by the results in [29]. A new equation, which is able to characterize the exact livenss probability of staying with the safe set $\mathcal{X}$, is formulated, and optimizations for addressing the safe probabilistic invariance problem are constructed via relaxing this equation.
@@ 2.2 Stochastic Continuous-time Systems
Based on the Doob’s nonnegative supermartingale inequality, barrier certificates were first proposed in [17] to address the safety verification problem in stochastic continuous-time systems. In [17], the safety verification problem aims to certify the lower bound of the probability that, starting from a given initial set of states, the system never enters a set of unsafe states when it evolves inside a specified state-constrained set. Subsequently, various barrier certificates were developed, each offering different levels of expressiveness and capabilities in handling diverse problems, e.g., inspired by the results in [12] and the Doob’s nonnegative supermartingale inequality, [22, 20, 21] proposed barrier-like conditions to tackle the finite-time safety verification problem; temporal logic verification of stochastic systems via barrier certificates was proposed in [9], which is to find a lower bound on the probability that a complex temporal logic property is satisfied by finite trajectories of the system. Recently, controlled versions, known as control barrier functions, were explored in [27, 14]. These controlled versions aim to facilitate the synthesis of controllers that guarantee the invariance of a specified safe set over finite or infinite time horizons, with a probability greater than a specified threshold. On the other hand, an alternate approach, distinct from the aforementioned nonnegative supermartingale based barrier certificates, was presented in [31]. It propose a system of equations capable of accurately describing the probability of the system eventually reaching a target set while remaining within a specified safe set before the first target hitting time. By relaxing this system of equations, it becomes possible to establish both a barrier-like condition for lower-bounding the probability and a barrier-like condition for upper-bounding the probability. Recently, this method was extended to lower- and upper-bound finite-time reachability probabilities in [28]. In this paper, we adapt the equation from [31] to precisely characterize the liveness probability and obtain a set of barrier-like conditions by relaxing the adapted equation for both lower- and upper-bounding the liveness probability for continuous-time systems.
@@
This paper is structured as follows. In Section 3, we formalize the stochastic discrete-time systems and associated safe probabilistic invariance verification of interest. In Section 4 we present two sets of optimizations designed to address the safe probabilistic invariance verification for stochastic discrete-time systems. In a parallel manner to Section 3, we extend our formalization to stochastic continuous-time systems in Section 5. Similarly, Section 6 is dedicated to presenting two sets of optimizations specifically designed for addressing safe probabilistic invariance verification for stochastic continuous-time systems. Examples demonstrating the proposed optimizations are provided in Section 7, and finally, this paper is concluded in Section 8.
@@ 3.1 Problem Statement
$\begin{cases}\bm{x}(l+1)=\bm{f}(\bm{x}(l),\bm{d}(l)),\quad\forall l\in\mathbb{N}\\ \bm{x}(0)=\bm{x}_{0}\in\mathbb{R}^{n}\end{cases}.$ (1)
@@
Given a disturbance signal $\pi$ and an initial state $\bm{x}_{0}\in\mathbb{R}^{n}$, a unique trajectory $\bm{\phi}_{\pi}^{\bm{x}_{0}}(\cdot)\colon\mathbb{N}\rightarrow\mathbb{R}^{n}$ is induced with $\bm{\phi}_{\pi}^{\bm{x}_{0}}(0)=\bm{x}_{0}$. Specifically, we have
$\bm{\phi}_{\pi}^{\bm{x}_{0}}(l+1)=\bm{f}(\bm{\phi}_{\pi}^{\bm{x}_{0}}(l),\bm{d}(l))$
for all $l\in\mathbb{N}$.
Given a safe set $\mathcal{X}$ and an initial set $\mathcal{X}_{0}$, where $\mathcal{X}_{0}\subseteq\mathcal{X}$, the safe probabilistic invariance verification problem is to determine lower and upper bounds of the liveness probability of remaining within the safe set $\mathcal{X}$ for system (1), starting from $\mathcal{X}_{0}$.
Definition 2
The safe probabilistic invariance verification for system (1) is to compute lower and upper bounds, denoted by $\epsilon_{1}\in[0,1]$ and $\epsilon_{2}\in[0,1]$ respectively, for the liveness probability that the system, starting from any state in $\mathcal{X}_{0}$, will remain inside the safe set $\mathcal{X}$ for all time, i.e., to compute $\epsilon_{1}$ and $\epsilon_{2}$ such that
$\epsilon_{1}\leq\mathbb{P}^{\infty}(\forall k\in\mathbb{N}.\bm{\phi}_{\pi}^{\bm{x}_{0}}(k)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0})\leq\epsilon_{2}.$ (2)
Remark 1
If the set $\mathcal{X}^{c}$ represents the desired (or, more comfortable) states, then $\epsilon_{2}$ can be used as an upper bound for the probability of the system (1) getting trapped within $\mathcal{X}$. This means that if we can ensure that the liveness probability is below $\epsilon_{2}$, we can be confident that the system will eventually reach the desired states with high probability.
We will focus on addressing the safe probabilistic invariance verification problem defined in Definition 2.
@@ 3.2 Reachability Probability Characterization in [29]
In this subsection, we will recall an equation that was derived for probabilistic reach-avoid analysis of stochastic discrete-time systems. The equation’s bounded solution is equivalent to the precise probability of the system entering a specified target set within a finite time while remaining inside a given safe set before the first target is reached. We will adopt this equation to address the safe invariance verification problem outlined in Definition 2.
Theorem 1 (Theorem 1, [29])
@@
$\begin{cases}v(\bm{x})=\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}^{\bm{x}}_{\pi}(1))]\\ v(\bm{x})=1_{\mathcal{X}_{r}}(\bm{x})+\mathbb{E}^{\infty}[w(\widehat{\bm{\phi}}^{\bm{x}}_{\pi}(1))]-w(\bm{x})\end{cases}$ (3)
then
$\begin{split}&\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\wedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(l)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}\\ &=\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\mid\bm{x}_{0}\in\mathcal{X}\big{)}\\ &=\lim_{i\rightarrow\infty}\frac{\mathbb{E}^{\infty}[\sum_{j=0}^{i-1}1_{\widehat{\mathcal{X}}\setminus\mathcal{X}}(\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(j))]}{i}=v(\bm{x}),\end{split}$
@@
A sufficient condition for certifying lower bounds of the probability $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\wedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(l)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}$ can be derived via relaxing (3). It is obtained by adding a constraint $v(\bm{x})\geq\epsilon_{1},\forall\bm{x}\in\mathcal{X}_{0}$ into the ones in Corollary 2 in [29].
Proposition 1
@@
$\begin{cases}v(\bm{x})\geq\epsilon_{1}&\forall\bm{x}\in\mathcal{X}_{0}\\ v(\bm{x})\leq\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]&\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}\\ v(\bm{x})\leq\mathbb{E}^{\infty}[w(\bm{\phi}^{\bm{x}}_{\pi}(1))]-w(\bm{x})&\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}\\ v(\bm{x})\leq 1&\forall\bm{x}\in\mathcal{X}_{r}\\ v(\bm{x})\leq 0&\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}\end{cases},$ (4)
then $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\wedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(l)\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}=\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}\geq\epsilon_{1}$.

### v2

@@ Abstract
Ensuring safety through set invariance has proven to be a valuable method in various robotics and control applications. This paper introduces a comprehensive framework for the safe probabilistic invariance verification of both discrete- and continuous-time stochastic dynamical systems over an infinite time horizon. The objective is to ascertain the lower and upper bounds of liveness probabilities for a given safe set and set of initial states. The liveness probability signifies the likelihood of the system remaining within the safe set indefinitely, starting from a state in the initial set. To address this problem, we propose optimizations for verifying safe probabilistic invariance in discrete-time and continuous-time stochastic dynamical systems. These optimizations are constructed via either using the Doob’s nonnegative supermartingale inequality-based method or relaxing the equations described in [30, 32], which can precisely characterize the probability of reaching a target set while avoiding unsafe states. Finally, we demonstrate the effectiveness of these optimizations through several examples using semi-definite programming tools.
@@
Stochastic Systems; Safe Probabilistic Invariance Verification; Liveness Probabilities; Lower and Upper Bounds.
@@ 1 Introduction
This paper proposes a framework for safe probabilistic invariance verification in both stochastic discrete-time and continuous-time systems. The objective of safe probabilistic invariance verification is to compute lower and upper bounds on liveness probabilities of the system remaining inside a safe set for all time, starting from a specific initial set. In this framework, several optimizations are proposed for computing these lower and upper bounds. These optimizations are constructed via the adaptation of stochastic barrier certificates in [3, 18], which are built upon the well-established Doob’s nonnegative supermartingale inequality [27], and equations in [30, 32], which can precisely characterize the probability of reaching a target set while avoiding unsafe states. To demonstrate and compare the performance of these optimizations, several examples are provided based on the semi-definite programming tool.
@@
    A framework for the safe probabilistic invariance verification in both stochastic discrete-time and continuous-time systems is proposed. The framework not only computes commonly studied lower bounds of the liveness probabilities, but also provides upper bounds. The incorporation of upper bounds enhances our understanding of the system’s invariance properties and allows for a more accurate estimation of the liveness probabilities.
@@
    The framework incorporates several optimizations to address the safe probabilistic invariance verification. These optimizations are derived from the known Doob’s nonnegative supermartingale inequality, as well as the equations proposed in [30, 32] for characterizing the exact probability. It is demonstrated that the optimizations derived from the nonnegative Doob’s supermartingale inequality are equivalent to those derived from the equations when computing lower bounds. However, for computing upper bounds, the optimizations from [30, 32] are more effective.
@@
This paper extends our previous conference work [33], which focused solely on the safe probabilistic invariance verification in stochastic discrete-time systems. In contrast to [33], this work expands the method to encompass continuous-time systems. Furthermore, we demonstrate that the optimizations derived from the nonnegative Doob’s supermartingale inequality are equivalent to those derived from the equations proposed by [30, 32] when computing lower bounds.
@@ 2.1 Stochastic Discrete-time Systems
The problem addressed in this work is related to the computation of probabilistic invariant sets. These sets define a range of states where a system, starting from any point within the set, must remain within a specified region of interest with a certain probability. By computing probabilistic invariant sets, we can demonstrate that the safe invariance property holds for an initial set if it is a subset of the computed probabilistic invariant sets. Previous works, such as [10], [11], and [8], have approximated polyhedral probabilistic invariant sets using Chebyshev’s inequality for linear systems with Gaussian noise. However, these methods are limited to computing lower bounds of the liveness probabilities and cannot be applied to compute upper bounds. Furthermore, the above mentioned works have primarily focused on linear systems, whereas this paper considers nonlinear systems.
@@
On the other hand, probabilistic invariance can be evaluated by examining its dual, probabilistic reachability. [2, 1] investigated the finite-time probabilistic invariance problem for discrete-time stochastic (hybrid) systems via reachability analysis. Meanwhile, [25, 26] studied the infinite-time probabilistic invariance by defining it as a finite-time reach-avoid property in combination with infinite-time invariance around absorbing sets over the state space and provided a lower bound for the infinite-time probabilistic invariance. Recently, [7] proposed an algorithm to compute infinite-horizon probabilistic controlled invariant sets which provide a lower bound on the liveness probabilities based on the dynamic program in [2]. These sets are computed using the stochastic backward reachable set from a robust invariant set. One obstacle faced by the aforementioned methods is the computation of absorbing or robust invariant sets, when they do exist.
@@
Another two closely related works to the present one are [3], which studies the safety and reachability verification based on non-negative supermartingale-based barrier certificates, and [30], which formulates a set of equations being able to characterize the exact probability of reaching a specified target set while avoiding unsafe states. With an assumption that the evolution space $\mathcal{X}$ is a robust invariant (i.e., $\bm{f}(\bm{x},\bm{d}):\mathcal{X}\times\mathcal{D}\rightarrow\mathcal{X}$), barrier certificates inspired by the ones in [18] and [19] were formulated for safety and reachability verification of stochastic discrete-time systems in [3]. The first set of optimizations for discrete-time systems in this paper was adapted from them. The assumption that the evolution space $\mathcal{X}$ (i.e., the safe set in this paper) is a robust invariant is abandoned in our method. Instead, an auxiliary switched system with a robust invariant set is borrowed to construct our constraints for addressing the safe probabilistic invariance verification problem. The second set of optimizations proposed in this paper is inspired by the results in [30]. A new equation, which is able to characterize the exact liveness probability of staying with the safe set $\mathcal{X}$, is formulated, and optimizations for addressing the safe probabilistic invariance verification problem are constructed via relaxing this equation.
@@ 2.2 Stochastic Continuous-time Systems
Based on the Doob’s nonnegative supermartingale inequality, barrier certificates were first proposed in [18] to address the safety verification problem in stochastic continuous-time systems. In [18], the safety verification problem aims to certify the lower bound of the probability that, starting from a given initial set of states, the system never enters a set of unsafe states when it evolves inside a specified state-constrained set. Subsequently, various barrier certificates were developed, each offering different levels of expressiveness and capabilities in handling diverse problems, e.g., inspired by the results in [13] and the Doob’s nonnegative supermartingale inequality, [23, 21, 22] proposed barrier-like conditions to tackle the finite-time safety verification problem; temporal logic verification of stochastic systems via barrier certificates was proposed in [9], which is to find a lower bound on the probability that a complex temporal logic property is satisfied by finite trajectories of the system. Recently, controlled versions, known as control barrier functions, were explored in [28, 15]. These controlled versions aim to facilitate the synthesis of controllers that guarantee the invariance of a specified safe set over finite or infinite time horizons, with a probability greater than a specified threshold. On the other hand, an alternate approach, distinct from the aforementioned ones based on the Doob’s nonnegative supermartingale inequality, was presented in [32]. It proposes a system of equations capable of accurately describing the probability of the system eventually reaching a target set while remaining within a specified safe set before the first target hitting time. By relaxing this system of equations, it becomes possible to establish both a barrier-like condition for lower-bounding the probability and a barrier-like condition for upper-bounding the probability. Recently, this method was extended to lower- and upper-bound finite-time reachability probabilities in [29]. In this paper, we adapt the equation from [32] to precisely characterize the liveness probability and obtain a set of barrier-like conditions by relaxing the adapted equation for both lower- and upper-bounding the liveness probabilities of continuous-time systems.
@@
This paper is structured as follows. In Section 3, we formalize stochastic discrete-time systems and associated safe probabilistic invariance verification problems of interest. In Section 4 we present two sets of optimizations designed to address the safe probabilistic invariance verification problem for stochastic discrete-time systems. In a parallel manner to Section 3, we extend our formalization to stochastic continuous-time systems in Section 5. Similarly, Section 6 is dedicated to presenting two sets of optimizations specifically designed for safe probabilistic invariance verification of stochastic continuous-time systems. Examples demonstrating the proposed optimizations are provided in Section 7, and finally, this paper is concluded in Section 8.
@@ 3.1 Problem Statement
(1) $\begin{cases}\bm{x}(l+1)=\bm{f}(\bm{x}(l),\bm{d}(l)),\quad\forall l\in\mathbb{N}\\ \bm{x}(0)=\bm{x}_{0}\in\mathbb{R}^{n}\end{cases}.$
@@
Given a disturbance signal $\pi$ and an initial state $\bm{x}_{0}\in\mathbb{R}^{n}$, a unique trajectory $\bm{\phi}_{\pi}^{\bm{x}_{0}}(\cdot)\colon\mathbb{N}\rightarrow\mathbb{R}^{n}$ is induced with $\bm{\phi}_{\pi}^{\bm{x}_{0}}(0)=\bm{x}_{0}$. Specifically, we have $\bm{\phi}_{\pi}^{\bm{x}_{0}}(l+1)=\bm{f}(\bm{\phi}_{\pi}^{\bm{x}_{0}}(l),\bm{d}(l))$ for all $l\in\mathbb{N}$.
Given a safe set $\mathcal{X}$ and an initial set $\mathcal{X}_{0}$, where $\mathcal{X}_{0}\subseteq\mathcal{X}$, the safe probabilistic invariance verification problem is to determine lower and upper bounds of liveness probabilities of remaining within the safe set $\mathcal{X}$ for system (1), starting from $\mathcal{X}_{0}$.
Definition 3.2.
The safe probabilistic invariance verification for system (1) is to compute lower and upper bounds, denoted by $\epsilon_{1}\in[0,1]$ and $\epsilon_{2}\in[0,1]$ respectively, for the liveness probabilities that the system, starting from $\mathcal{X}_{0}$, will remain inside the safe set $\mathcal{X}$ for all time, i.e., to compute $\epsilon_{1}$ and $\epsilon_{2}$ such that
(2) $\epsilon_{1}\leq\mathbb{P}^{\infty}(\forall k\in\mathbb{N}.\bm{\phi}_{\pi}^{\bm{x}_{0}}(k)\in\mathcal{X})\leq\epsilon_{2},\forall\bm{x}_{0}\in\mathcal{X}_{0}.$
Remark 3.3.
If the set $\mathcal{X}^{c}$ represents the desired (or, more comfortable) states, then $\epsilon_{2}$ can be used as an upper bound for the probabilities of the system (1) getting trapped within $\mathcal{X}$. This means that if we can ensure that all the liveness probabilities are below $\epsilon_{2}$, we can be confident that the system will eventually reach the desired states with high probability.
We will focus on addressing the safe probabilistic invariance verification problem defined in Definition 3.2.
@@ 3.2 Reachability Probability Characterization in [30]
In this subsection, we will recall an equation that was derived for probabilistic reach-avoid analysis of stochastic discrete-time systems. The equation’s bounded solution is equivalent to the precise probability of the system entering a specified target set within a finite time while remaining inside a given safe set before the first target is reached. We will adopt this equation to address the safe invariance verification problem outlined in Definition 3.2.
Proposition 3.4 (Theorem 1, [30]).
@@
(3) $\begin{cases}v(\bm{x})=\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}^{\bm{x}}_{\pi}(1))]\\ v(\bm{x})=1_{\mathcal{X}_{r}}(\bm{x})+\mathbb{E}^{\infty}[w(\widehat{\bm{\phi}}^{\bm{x}}_{\pi}(1))]-w(\bm{x})\end{cases}$
then for $\bm{x}_{0}\in\mathcal{X}$,
$\begin{split}&\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\wedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(l)\in\mathcal{X}\big{)}\\ &=\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\big{)}=\lim_{i\rightarrow\infty}\frac{\mathbb{E}^{\infty}[\sum_{j=0}^{i-1}1_{\widehat{\mathcal{X}}\setminus\mathcal{X}}(\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(j))]}{i}=v(\bm{x}_{0})\end{split},$
@@
A sufficient condition for certifying lower bounds of the probabilities $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\wedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(l)\in\mathcal{X}\big{)}$ for $\bm{x}_{0}\in\mathcal{X}_{0}$ can be derived via relaxing (3). It is obtained by adding a constraint $v(\bm{x})\geq\epsilon_{1},\forall\bm{x}\in\mathcal{X}_{0}$ into the ones in Corollary 2 in [30].
Corollary 3.5.
@@
(4) $\begin{cases}v(\bm{x})\geq\epsilon_{1}&\forall\bm{x}\in\mathcal{X}_{0}\\ v(\bm{x})\leq\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]&\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}\\ v(\bm{x})\leq\mathbb{E}^{\infty}[w(\bm{\phi}^{\bm{x}}_{\pi}(1))]-w(\bm{x})&\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}\\ v(\bm{x})\leq 1&\forall\bm{x}\in\mathcal{X}_{r}\\ v(\bm{x})\leq 0&\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}\end{cases},$
then $\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\wedge\forall l\in\mathbb{N}_{\leq k}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(l)\in\mathcal{X}\big{)}=\mathbb{P}^{\infty}\big{(}\exists k\in\mathbb{N}.\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X}_{r}\big{)}\geq\epsilon_{1}$ for $\bm{x}_{0}\in\mathcal{X}_{0}$.